OpenStudy (king2001):

Using a reference triangle, find the exact value of the expression: cos(tan^-1 x^2) A.1/sqrt x^4+1 B.1/x^4-1 C.x^4/sqrt x^4+1 D.x^4/x^4+1 I answered A

OpenStudy (thewafflebro):

ok noob lemme see your problem

OpenStudy (king2001):

thats my problem above i answered A

OpenStudy (john_es):

I think you mean, \[\cos(\arctan(x^2))\] right?

OpenStudy (king2001):

yes

OpenStudy (john_es):

Well, there is an intesting formula that relates the cos and the arctan in this case. The question \[\cos(\arctan(x^2))\] means: "calculate the cos of the angle whose tangent is equals to x^2"

OpenStudy (john_es):

That means, \[\tan\alpha=x^2\]And you know \[1+\tan^2\alpha=\frac{1}{\cos^2\alpha}\]So, \[\cos\alpha=\frac{1}{\sqrt{1+\tan^2\alpha}}\]That means, \[\cos\alpha=\frac{1}{\sqrt{1+x^2}}\]

OpenStudy (king2001):

yes

OpenStudy (john_es):

Sorry, the last formula should be, \[\cos\alpha=\frac{1}{\sqrt{1+x^4}}\]And this is the answer.

OpenStudy (john_es):

Well, I tried to do the process, but it seems I also give the answer.

OpenStudy (king2001):

ok thanks

OpenStudy (john_es):

You're welcome.

OpenStudy (king2001):

i have a couple more problems i need help with

OpenStudy (john_es):

Tell me.

OpenStudy (king2001):

here or on the forum

OpenStudy (john_es):

Wherever you want.

OpenStudy (king2001):

ok ill tell here

OpenStudy (king2001):

Identify the middle line for the function: f(t)=1/2sec(t+2)-pi A.y=-pi B.y=pi C.y=pi/2 D.y=-pi/2 I answered C

OpenStudy (john_es):

Let me see.

OpenStudy (john_es):

You mean the function, \[f(t)=\frac{1}{2\sec(t+2)}-\pi\]?

OpenStudy (king2001):

no

OpenStudy (john_es):

Ok, I see now.

OpenStudy (john_es):

Well, the midline of a trigonometric function I think is another point.

OpenStudy (king2001):

yes

OpenStudy (john_es):

Do you remember the graph of the sine? The midline I think is the line where you measure the amplitude. Or better is the line where the sine is zero.

OpenStudy (john_es):

In this case, is the line where the function with t is zero

OpenStudy (john_es):

So the midline is in -pi

OpenStudy (king2001):

ok

OpenStudy (john_es):

Just the first answer.

OpenStudy (king2001):

so the answer is A?

OpenStudy (john_es):

Yes, the line where the function with t is zero.

OpenStudy (king2001):

ok i have i think 1 more problem

OpenStudy (king2001):

i answered B

OpenStudy (john_es):

Perfect!

OpenStudy (king2001):

ok good thanks

OpenStudy (john_es):

You're welcome.

OpenStudy (king2001):

ok 1 more

OpenStudy (king2001):

i answered A is that also correct or i did it wrong

OpenStudy (king2001):

john?

OpenStudy (john_es):

Ok, you're answer is perfect!

OpenStudy (king2001):

wow really?

OpenStudy (king2001):

ok last problem i promise

OpenStudy (john_es):

Yes, the number that multiplies the t gives you the angular frequency, \[2\pi f=1\Rightarrow f=1/(2\pi)\] The period is the inverse of the frequency, so, \[T=1/f=2\pi\]

OpenStudy (king2001):

ok what about this last problem i just sent

OpenStudy (john_es):

Well, what is the value of the y if the function with t goes to zero?

OpenStudy (king2001):

zero?

OpenStudy (john_es):

Exactly.

OpenStudy (king2001):

so would it be y=0?

OpenStudy (john_es):

Yes :)

OpenStudy (king2001):

ok thanks so much!

OpenStudy (john_es):

You're welcome.